OpenStudy (anonymous):

The expression ½logm - 3logn is equivalent to 1) log√m+logn^3 3) log ((m^2)/(3√n) 2) log½m + 3logn 4) log (√ m)/n^3

OpenStudy (anonymous):

it's equivalent to 4) using the log propreties a*log(x)=log(x^a) and log(a/b)=log(a)-log(b)